OpenStudy (anonymous):

A recipe for potting soil that makes enough to fill 4 bags calls for 6 1/4 lb of compost. Kailee wants to modify the recipe so that it fills 10 bags. How many more pounds of compost will Kailee need for the modified recipe than the original recipe calls for?

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

A. 2 17/32

OpenStudy (anonymous):

B. 9 3/8

OpenStudy (anonymous):

C 7 13/16

OpenStudy (anonymous):

15 5/8

OpenStudy (anonymous):

that was d

OpenStudy (anonymous):

i ithink its c

OpenStudy (anonymous):

help @Jennaluv

OpenStudy (anonymous):

hhelp me @mathmath333

OpenStudy (anonymous):

@HelpOfTheGods

OpenStudy (mathmath333):

\(\Large\rm \dfrac{61}{4} ib=4compost\) \(\Large\rm \dfrac{61}{4} \times \dfrac{10}{4}ib=4\times \dfrac{10}{4}compost\) \(\Large \rm Answer = \dfrac{61}{4} \times \dfrac{10}{4}ib\)

OpenStudy (anonymous):

not ts not 61/4 its 6 1/4

OpenStudy (mathmath333):

oh

OpenStudy (anonymous):

6 4 ____ 1

OpenStudy (mathmath333):

u mean 25/4 right

OpenStudy (anonymous):

/for the math problem no

OpenStudy (mathmath333):



OpenStudy (mathmath333):

u can draw and tell me exact thing

OpenStudy (anonymous):

draw what!!

OpenStudy (mathmath333):

draw 6 1/4

OpenStudy (anonymous):